OpenStudy (anonymous):

Enrollment went up 625 to 710 over a year what is the percent of increase?

OpenStudy (shinalcantara):

First, get the increase 710 - 625 = ?

OpenStudy (shinalcantara):

@mmehope

OpenStudy (shinalcantara):

after getting the increase, divide it by the original number which is 625 to get it's decimal equivalent. then multiply by 100% and you'll have it in percentage

OpenStudy (shinalcantara):

\[\frac{ 710-625 }{ 625 } \times 100% = ?\]

OpenStudy (shinalcantara):

Simplify and you'll get the percentage

OpenStudy (anonymous):

Ok thanks, any advice on problems like this in the future? How to know when to use this set up?